### v6

@@ 2.3 Analysis of joint pretraining
To the best of our knowledge, BrainWave is the first foundational model that combines invasive and non-invasive neural data. We next examine the effectiveness of this joint pretraining strategy to determine whether it is more effective to pretrain a separate model for each recording type and apply it for corresponding downstream tasks with the same recording type, or to utilize a joint pretraining approach. To this end, we separately pretrained two model variants, namely BrainWave-EEG and BrainWave-iEEG, using EEG and iEEG data, respectively, while keeping the model and pretraining configurations (Supplementary Tables 19 and 18) identical to BrainWave. The numbers of patches for pretraining BrainWave-EEG and BrainWave-iEEG are relatively balanced (1.74 billion versus 1.42 billion). Subsequently, we conducted cross-domain evaluation and few-shot classification on both model variants, in which BrainWave-EEG was evaluated on EEG datasets and BrainWave-iEEG was evaluated on iEEG datasets.
@@
Through a series of experiments, we observed that BrainWave outperformed the other two variants in almost all tasks and experimental settings (Fig. 4a-d and Extended Data and Figs. 12 and 14), except in experiment Absence-16 to Atonic-5 where BrainWave is slightly lower than BrainWave-EEG in terms of BACC (44.55% versus 47.23%). From the overall results, we discovered that the integration of iEEG during joint pretraining enhanced the performance on downstream tasks based on EEG data and vice versa, suggesting that BrainWave is able to capture fundamental insights about brain activities by combining two distinct types of signals. In comparison to the average improvement over two model variants (Fig. 4e), the improvement of the BrainWave over BrainWave-EEG was more significant than the improvement over BrainWave-iEEG (7.44% versus 2.58% on cross-domain evaluation and 8.13% versus 6.92% on few-shot classification in terms of AUROC). The phenomenon suggested that the boost in performance achieved by incorporating iEEG data in the pretraining was more pronounced, which might be attributed to the lower signal-to-noise ratio and higher accuracy of intracranial neural signals.
@@
Given that joint pretraining leads to increase in performance, we further explored the underlying reasons by analyzing the representations learned by the models. First, we validated whether joint pretraining can learn more enriched information. For this purpose, we performed principal component analysis (PCA) to the pretrained representations, selecting principal components until 99% of the variance could be explained. We conducted analysis on 15 datasets and recorded the number of principal components $k$ of each dataset. On 13 out 14 datasets, BrainWave yielded a higher $k$ value (Supplementary Tables 20 and 21), indicating that BrainWave demonstrated the ability to extract more enriched information from downstream datasets. Furthermore, we conducted an additional experiment by evaluating the performance of the BrainWave versus other variants on another type of biosignal, electrocardiogram (ECG), through which we aimed to verify if joint pretraining results in a stronger adaptability on new tasks due to the acquisition of more general patterns. The results (Fig. 4f) showed that BrainWave achieves an improvement of 9.90% and 19.48% in terms of AUROC over BrainWave-iEEG and BrainWave-EEG, respectively, demonstrating that joint pretraining enables better generalization to unseen data types. To summarize, BrainWave learned richer semantic information and more general patterns of the data than other model variants with only one type of data. This finding opens up possibilities for expanding signal types and developing more versatile foundational models for biosignals.
@@ 3 Discussion
We have introduced BrainWave, a brain signal foundation model that learns robust representations of electrical brain recordings for a broad range of clinical applications. To the best of our knowledge, BrainWave is the first model pretrained on a large-scale dataset composed of recordings from both invasive and non-invasive methods, which comprised more than 3 billion signal patches from approximately 16,000 individuals. We adopted a masked modeling strategy to pretrain BrainWave, in which the model learned to reconstruct the whole sequence given its partial observation. The architecture design allowed for brain recordings of varying lengths, sampling rates, and electrode counts, enhancing the flexibility of BrainWave for joint pretraining and deployment on iEEG and EEG data. In comprehensive experiments involving cross-domain evaluation and few-shot classification, we demonstrated the versatility of BrainWave across a wide array of healthcare tasks under different settings, which indicated the great potential of BrainWave assist clinical diagnostics and health condition detection.
Beyond its practical value, BrainWave also provides valuable insights for the development of foundation models based on brain signals. As the first study to adopt joint pretraining with EEG and iEEG data, we validated the effectiveness of this strategy. Through extensive evaluation, we observed a clear performance boost from joint pretraining compared to the model variants from control group. Furthermore, we revealed the reasons behind the performance enhancement attributed to joint training. Our work opens up new opportunities for leveraging similar techniques with other types of medical data, where efficient and scalable models are critical. In addition, this study provides a foundation for further exploration into joint pretraining across multiple medical data modalities. Future studies can build upon our findings to investigate whether similar improvements can be achieved with other neural data types or even cross-domain data involving different physiological signals.
@@
Despite the promising results, there is a wealth of potential for further development and advancement. Firstly, BrainWave is a unimodal model and cannot handle data from other modalities, such as magnetic resonance imaging (MRI), which can provide higher spatial resolution compared to electrical signals and is widely used in healthcare applications. Therefore, we hope to design a model framework going forward that is compatible with multiple modalities. Secondly, diverse medical scenarios collect various physiological signals, with diagnoses sometimes relying on multiple signal types. For instance, stroke diagnosis and rehabilitation often require the recording of EEG and EMG (Electromyography)[54]. To better serve more healthcare applications, the model should be capable of handling a wider range of biosignals. In this study, we only made preliminary attempts using ECG data, suggesting that there is still progress to be made before a comprehensive model capable of handling a diverse range of physiological signals is developed.
@@ 4.1 Technique details
Fig. 1b shows the overall architecture of BrainWave, which is composed of three main components: embedding layer, Transformer encoder and channel attention. This section aims to introduce the details of these components.
Embedding layer. The embedding layer of BrainWave projects the original signals into latent embeddings, in which each channel is operated independently. Given a single-channel brain recording, we first divided the signal into a series of consecutive non-overlapping 1-second patches $\mathbf{P}_{i}\in\mathbb{R}^{N\times P}$, where $N$ is the number of patches, $P$ is the number of timestamps in each 1-second patch, and $i$ is the channel index. Then we calculated the time-frequency representations of each patch, in which we chose the spectrograms with Gaussian window. By keeping the ratio of the window size and the hop size to the patch length $P$ constant, we can align recordings with different sampling rates onto spectrograms with consistent temporal and frequency resolutions. Specifically, we set the window size equal to $\frac{P}{4}$ and the hop size equal to $\frac{P}{8}$, and the resulting spectrograms are denoted as $\mathbf{S}_{i}\in\mathbb{R}^{N\times T\times F}$, where $T$ is the length of the time axis and $F$ is the length of the frequency axis. We convolved $\mathbf{S}_{i}$ using 2D convolutional kernels to obtain the feature maps $\mathbf{M}_{i}\in\mathbb{R}^{N\times C_{\text{out}}\times T_{\text{out}}\times F_{\text{out}}}$, where $C_{\text{out}}$ is the output channels and $T_{\text{out}}\times F_{\text{out}}$ is the size of the feature maps. Since our design ensured that data with different sampling rates have the same temporal and frequency resolution, the length of the time axis $T_{\text{out}}$ in the resulting feature maps was the same (because the duration of all patches is 1 second), while the length of the frequency axis $F_{\text{out}}$ varied. Therefore, we performed padding or truncation on the frequency axis to standardize the size of the feature maps. Then we flattened the standardized feature maps and projected them with a linear layer to derive the input embeddings $\mathbf{E}_{i}\in\mathbb{R}^{N\times D}$, where $D$ is the hidden size.
@@ 4.2 Pretraining
The preprocessing of the pretraining data primarily involved channel selection, downsampling and filtering. Due to potential equipment issues during the data collection process, there might be invalid channels where no valid brain signals were captured. Therefore, we needed to perform channel selection, in which we visualized the recordings and manually selected the valid dchannels. We only downsampled signals with a sampling rate higher than 1000 Hz until they were below 1000 Hz, with the remaining signals preserved at their original sampling rates. Then, we applied a bandpass filter in the frequency range of 0.01 Hz to sfreq/3 Hz, where sfreq represents the sampling rate, and applied a notch filter at 50 Hz or 60 Hz to remove powerline noise.
@@ 4.3 Downstream evaluation
Evaluation datasets. The evaluation benchmark comprised 15 distinct datasets, which contain 7 different healthcare scenarios.
@@
Epilepsy: The CHB-MIT [49, 68] database consists of EEG recordings from 22 pediatric subjects (5 males, ages 3–22; and 17 females, ages 1.5–19) with intractable seizures. We randomly split the subjects into 5 groups. The data comprised 23 channels with a sampling rate of 256 Hz. We split the data into 10-second segments and obtained 4148 samples in total. Absence-16, Clonic-6 and Atonic-5 are 3 private datasets that were collected from patients with absence seizures, clonic seizures and atonic seizures, respectively. The annotations were divided into three categories: epileptic waveforms, normal waveforms, and Interictal epileptiform discharge (IED). The recordings comprised 19 channels with a sampling rate of 256 Hz. After processing the data into 4-second segments, we obtained a total of 8016 samples from Absence-16, 2426 samples from Clonic-6, and 1587 samples from Atonic-5. We randomly divided the 16 patients in Absence-16 into 5 groups. The Mayo-Clinic [52] data were collected between 1 AM and 3 AM from 25 patients with drug resistant epilepsy (DRE) undergoing evaluation for epilepsy surgery. The FNUSA [52] dataset is made up of iEEG data collected in awake resting state from 14 patients diagnosed with DRE. These patients underwent a standard pre-surgical monitoring for localization of seizure onset zone (SOZ), a standard procedure for epilepsy surgery. We split Mayo-Clinic and FNUSA into 6 and 5 subject groups, respectively. Both Mayo-Clinic and FNUSA were segmented into 3-second data clips and downsampled to 1000 Hz. In order to locate the SOZ, annotations were made for each channel. We preserved the data segments annotated with physiological activity, pathological (epileptic) activity and artifacts. In total, Mayo-Clinic contained 113,260 samples and FNUSA contained 179,629 samples. DRE-Clinical is a private dataset that contains the iEEG recordings from 8 patients with DRE. The annotation was also channel-wise for SOZ localization. We split the patients into 4 groups, with two individuals in each group. The recordings were segmented into 20-second data clips and we obtained 229,932 samples in total.
Depression: MDD-64 [53] contains EEG recordings from 64 subjects with 34 of them diagnosed with Major Depressive Disorder (MDD). We randomly split the subjects into 5 groups. The data comprised 19 channels with a sampling rate of 256 Hz. We split the data into 10-second segments and obtained 7309 samples in total. Depression-122 [69] consists of resting EEG data with 122 college-age participants (47 males, ages 18–24; 74 females, ages 18–23; and 1 unknown) with their scores in Beck Depression Inventory (BDI). According to  [70], participants with BDI scores $>13$ were considered depressed. Healthy controls had stable low BDI scores ($<7$) and no self-reported history or symptoms of anxiety disorder. The data comprised 64 channels with a sampling rate of 500 Hz. We split the data into 10-second segments and obtained 5836 samples in total.
@@
Sleep Deprivation: SD-71 [71] provides resting-state EEG data from 71 participants who underwent two experiments involving normal sleep and sleep deprivation. We randomly split the subjects into 5 groups. The data comprised 61 channels with a sampling rate of 500 Hz. We split the data into 5-second segments and obtained 13,700 samples in total.
@@
Cross-subject evaluation. The cross-subject evaluation involved experiments on 12 datasets: AD-65 (Alzheimer’s disease diagnosis), CHB-MIT (seizure detection), Absence-16 (seizure detection), Mayo-Clinic (seizure detection and SOZ localization), FNUSA (seizure detection and SOZ localization), DRE-Clinical (seizure detection and SOZ localization), MDD-64 (MDD diagnosis), Depression-122 (depression diagnosis), Schizophrenia-28 (schizophrenia diagnosis), ADHD-Adult (ADHD diagnosis), ADHD-child (ADHD diagnosis) and SD-71 (sleep deprivation detection). We used the AdamW optimizer for fine-tuning, with $\beta_{1}=0.9$, $\beta_{2}=0.95$, $eps=10^{-5}$. For all the models, we fine-tuned the pretrained encoder and the classification head with a learning rate of $1\times 10^{-5}$ and $1\times 10^{-4}$. The models were trained for up to 30 epochs, and then the best-performing models on the validation set were selected for testing.
Cross-hospital and cross-subtype evaluation. The cross-hospital evaluation involved Mayo-Clinic and FNUSA, and the cross-subtype evaluation involved Absence-16, Clonic-6, and Atonic-5. In the experiments, we fine-tune the models on the source dataset with a fixed number of epochs and directly evaluated on the target dataset. We used the AdamW optimizer, with $\beta_{1}=0.9$, $\beta_{2}=0.95$, $eps=10^{-5}$. We fine-tuned the pretrained encoder and the classification head for 5 epochs with a learning rate of $1\times 10^{-5}$ and $1\times 10^{-4}$.

### v7

@@ 2.4 Analysis of Joint Pretraining
To the best of our knowledge, BrainWave is the first foundational model that combines invasive and non-invasive neural data. We next examine the effectiveness of this joint pretraining strategy to determine whether it is more effective to pretrain a separate model for each recording type and apply it for corresponding downstream tasks with the same recording type, or to utilize a joint pretraining approach. To this end, we separately pretrained two model variants, namely BrainWave-EEG and BrainWave-iEEG, using EEG and iEEG data, respectively, while keeping the model and pretraining configurations (Supplementary Tables 18 and 19) identical to BrainWave. The numbers of patches for pretraining BrainWave-EEG and BrainWave-iEEG are relatively balanced (1.74 billion versus 1.42 billion). Subsequently, we conducted cross-domain evaluation and few-shot classification on both model variants, in which BrainWave-EEG was evaluated on EEG datasets and BrainWave-iEEG was evaluated on iEEG datasets.
@@
Through a series of experiments, we observed that BrainWave outperformed the other two variants in almost all tasks and experimental settings (Fig. 5a-d and Extended Data and Figs. 11 and 12), except in experiment Absence-16 to Atonic-5 where BrainWave is slightly lower than BrainWave-EEG in terms of BACC (44.55% versus 47.23%). From the overall results, we discovered that the integration of iEEG during joint pretraining enhanced the performance on downstream tasks based on EEG data and vice versa, suggesting that BrainWave is able to capture fundamental insights about brain activities by combining two distinct types of signals. In comparison to the average improvement over two model variants (Fig. 5e), the improvement of the BrainWave over BrainWave-EEG was more significant than the improvement over BrainWave-iEEG (7.25% versus 2.05% on cross-domain evaluation and 8.30% versus 4.98% on few-shot classification in terms of AUROC). The phenomenon suggested that the boost in performance achieved by incorporating iEEG data in the pretraining was more pronounced, which might be attributed to the lower signal-to-noise ratio and higher accuracy of intracranial neural signals.
@@
Given that joint pretraining leads to increase in performance, we further explored the underlying reasons by analyzing the representations learned by the models. First, we validated whether joint pretraining can learn more enriched information. For this purpose, we performed principal component analysis (PCA) to the pretrained representations, selecting principal components until 99% of the variance could be explained. We conducted analysis on 12 datasets and recorded the number of principal components $k$ of each dataset. On 11 out 12 datasets, BrainWave yielded a higher $k$ value (Supplementary Tables 20 and 21), indicating that BrainWave demonstrated the ability to extract more enriched information from downstream datasets. Furthermore, we conducted an additional experiment by evaluating the performance of the BrainWave versus other variants on another type of biosignal, electrocardiogram (ECG), through which we aimed to verify if joint pretraining results in a stronger adaptability on new tasks due to the acquisition of more general patterns. The results (Fig. 5f) showed that BrainWave achieves an improvement of 9.90% and 19.48% in terms of AUROC over BrainWave-iEEG and BrainWave-EEG, respectively, demonstrating that joint pretraining enables better generalization to unseen data types. To summarize, BrainWave learned richer semantic information and more general patterns of the data than other model variants with only one type of data. This finding opens up possibilities for expanding signal types and developing more versatile foundational models for biosignals.
@@ 3 Discussion
We have introduced BrainWave, a brain signal foundation model that learns robust representations of electrical brain recordings for a broad range of clinical applications. To the best of our knowledge, BrainWave is the first model pretrained on a large-scale dataset composed of recordings from both invasive and non-invasive modalities, which comprised more than 3 billion signal patches from approximately 16,000 individuals. The architectural design accommodated brain recordings of varying lengths, sampling rates, and electrode counts, thereby enhancing the flexibility of BrainWave for joint pretraining and deployment on EEG and iEEG data. We employed a masked modeling strategy to pretrain BrainWave, enabling the model to reconstruct the complete sequence from partial observations. In comprehensive experiments involving cross-domain evaluation and few-shot classification, we demonstrated the versatility of BrainWave across a wide array of brain disorder detection tasks under diverse conditions. We also applied BrainWave to real-world clinical scenarios, suggesting that BrainWave holds significant potential to support clinical diagnostics and the identification of health conditions.
Beyond its practical value, BrainWave yields valuable insights into the fundamental model governing multimodal neural electrical signals. The present study represents the initial investigation to utilize joint pretraining with both EEG and iEEG data. Through rigorous assessment, we demonstrated the effectiveness of this novel methodological approach. The results of our study demonstrate a significant performance boost derived from the joint pretraining approach, compared to control model configurations. Additionally, we have elucidated the factors contributing to the enhanced performance associated with joint training. Future investigations can build upon our findings to examine whether comparable performance enhancements can be attained with other neural data types or even cross-domain data encompassing varied physiological signals.
As an exciting interdisciplinary research direction between neuroscience and artificial intelligence, BrainWave may also provide a new approach to deciphering the mechanisms of brain information processing. A key challenge in neuroscience research involves analyzing and elucidating the fundamental operating principles governing population-level neural networks, as derived from high-volume neural recording data. Large language models have demonstrated its capacity to extract fundamental attributes of human cognition by compressing substantial human linguistic data LABEL:yin2024entropylawstorydata,_naveed2024comprehensiveoverviewlargelanguage. Similarly, the impressive performance of BrainWave, as shown in the present study, may be attributed to their ability to thoroughly comprehend and effectively extract features of brain neural activity. Analyzing the network properties of BrainWave could potentially offer insights to guide research on biological neural networks. Specifically, by modulating various parameters of BrainWave, analogous models representing diverse brain disorders can be generated, thereby establishing a novel experimental framework for neuroscience and brain disease research.
@@
Despite the promising results, there is a wealth of potential for further development and advancement. Firstly, BrainWave cannot handle data from other modalities, such as magnetic resonance imaging (MRI), which can provide higher spatial resolution compared to electrical signals and is widely used in healthcare applications. Consequently, our ultimate objective is to develop a model framework that can accommodate various data modalities. Secondly, diverse medical scenarios collect various physiological signals, with diagnoses sometimes relying on multiple signal types. For instance, stroke diagnosis and rehabilitation often require the recording of EEG and EMG (Electromyography)[54]. To enhance suitability for a more expansive set of healthcare applications, the model should possess the capability to accommodate a more diverse array of biosignals. This investigation involved initial efforts utilizing electrocardiogram data, which indicates that additional advancements are necessary before constructing a comprehensive model able to accommodate a variety of physiological measurements.
@@ 4.1 Technique details
Fig. 1b shows the overall architecture of BrainWave, which is composed of three main components: scale alignment layer, Transformer encoder and channel attention. This section aims to introduce the details of these components.
Scale alignment layer. One of the largest challenges in modeling brain signals by a unified model is the diversity of sampling rates, which leads to inconsistent temporal resolutions. Previous works often address this issue by resampling all the signals into a common frequency. However, such a strategy becomes ineffective when dealing with both iEEG and EEG data. As sampling rates of iEEG and EEG differ greatly, resampling them to a shared scale is not only inflexible but may lead to a loss of signal fidelity. To overcome this problem, we propose a novel embedding approach which maps signals with arbitrary sampling rates into a space with a unified scale.
The scale alignment layer of BrainWave projects the original signals into latent embeddings, in which each channel is operated independently. Given a single-channel brain recording, we first divided the signal into a series of consecutive non-overlapping 1-second patches $\mathbf{P}_{i}\in\mathbb{R}^{N\times P}$, where $N$ is the number of patches, $P$ is the number of timestamps in each 1-second patch, and $i$ is the channel index. Then we calculated the time-frequency representations of each patch, in which we chose the spectrograms with Gaussian window. By keeping the ratio of the window size and the hop size to the patch length $P$ constant, we can align recordings with different sampling rates onto spectrograms with consistent temporal and frequency resolutions. Specifically, we set the window size equal to $\frac{P}{4}$ and the hop size equal to $\frac{P}{8}$, and the resulting spectrograms are denoted as $\mathbf{S}_{i}\in\mathbb{R}^{N\times T\times F}$, where $T$ is the length of the time axis and $F$ is the length of the frequency axis. We convolved $\mathbf{S}_{i}$ using 2D convolutional kernels to obtain the feature maps $\mathbf{M}_{i}\in\mathbb{R}^{N\times C_{\text{out}}\times T_{\text{out}}\times F_{\text{out}}}$, where $C_{\text{out}}$ is the output channels and $T_{\text{out}}\times F_{\text{out}}$ is the size of the feature maps. Since our design ensured that data with different sampling rates have the same temporal and frequency resolution, the length of the time axis $T_{\text{out}}$ in the resulting feature maps was the same (because the duration of all patches is 1 second), while the length of the frequency axis $F_{\text{out}}$ varied. Therefore, we performed padding or truncation on the frequency axis to standardize the size of the feature maps. Then we flattened the standardized feature maps and projected them with a linear layer to derive the input embeddings $\mathbf{E}_{i}\in\mathbb{R}^{N\times D}$, where $D$ is the hidden size.
@@ 4.2 Pretraining
The preprocessing of the pretraining data primarily involved channel selection and filtering. Due to potential equipment issues during the data collection process, there might be invalid channels where no valid brain signals were captured. Therefore, we needed to perform channel selection, in which we visualized the recordings and manually selected the valid channels. Because data acquisition may be affected by power line interference, we apply a 50 Hz or 60 Hz notch filter to remove power line noise. Since the AC power grid frequency varies across countries and regions (commonly 50 Hz or 60 Hz), the notch filter frequency is determined by the local power line frequency where the data is collected.
@@ 4.3 Downstream evaluation
Evaluation datasets. The evaluation benchmark comprised 13 distinct datasets.
@@
Epilepsy: The CHB-MIT [49, 68] database consists of EEG recordings from 22 pediatric subjects (5 males, ages 3–22; and 17 females, ages 1.5–19) with intractable seizures. We randomly split the subjects into 5 groups. The data comprised 23 channels with a sampling rate of 256 Hz. We split the data into 10-second segments and obtained 4148 samples in total. Absence-16, Clonic-6 and Atonic-5 are 3 private datasets that were collected from patients with absence seizures, clonic seizures and atonic seizures, respectively. The annotations were divided into three categories: epileptic waveforms, normal waveforms, and Interictal epileptiform discharge (IED). The recordings comprised 19 channels with a sampling rate of 256 Hz. After processing the data into 4-second segments, we obtained a total of 8016 samples from Absence-16, 2426 samples from Clonic-6, and 1587 samples from Atonic-5. We randomly divided the 16 patients in Absence-16 into 5 groups. The Mayo-Clinic [52] data were collected between 1 AM and 3 AM from 25 patients with DRE undergoing evaluation for epilepsy surgery. The FNUSA [52] dataset is made up of iEEG data collected in awake resting state from 14 patients diagnosed with DRE. We split Mayo-Clinic and FNUSA into 6 and 5 subject groups, respectively. Both Mayo-Clinic and FNUSA were segmented into 3-second data clips and downsampled to 1000 Hz. In order to locate the SOZ, annotations were made for each channel. We preserved the data segments annotated with physiological activity, pathological (epileptic) activity and artifacts. In total, Mayo-Clinic contained 113,260 samples and FNUSA contained 179,629 samples.
Depression: MDD-64 [53] contains EEG recordings from 64 subjects with 34 of them diagnosed with Major Depressive Disorder (MDD). We randomly split the subjects into 5 groups. The data comprised 19 channels with a sampling rate of 256 Hz. We split the data into 10-second segments and obtained 7309 samples in total. Depression-122 [69] consists of resting EEG data with 122 college-age participants (47 males, ages 18–24; 74 females, ages 18–23; and 1 unknown) with their scores in Beck Depression Inventory (BDI). According to  [70], participants with BDI scores >13>13 were considered depressed. Healthy controls had stable low BDI scores (<7<7) and no self-reported history or symptoms of anxiety disorder. The data comprised 64 channels with a sampling rate of 500 Hz. We split the data into 10-second segments and obtained 5836 samples in total.
@@
Cross-subject evaluation. The cross-subject evaluation involved experiments on 10 datasets: AD-65 (Alzheimer’s disease diagnosis), CHB-MIT (seizure detection), Absence-16 (seizure detection), Mayo-Clinic (seizure detection), FNUSA (seizure detection), MDD-64 (MDD diagnosis), Depression-122 (depression diagnosis), Schizophrenia-28 (schizophrenia diagnosis), ADHD-Adult (ADHD diagnosis) and ADHD-child (ADHD diagnosis). We used the AdamW optimizer for fine-tuning, with $\beta_{1}=0.9$, $\beta_{2}=0.95$, $eps=10^{-5}$. For all the models, we fine-tuned the pretrained encoder and the classification head with a learning rate of $1\times 10^{-5}$ and $1\times 10^{-4}$. The models were trained for up to 30 epochs, and then the best-performing models on the validation set were selected for testing.
Cross-hospital and cross-subtype evaluation. The cross-hospital evaluation involved Mayo-Clinic and FNUSA, and the cross-subtype evaluation involved Absence-16, Clonic-6, and Atonic-5. In the experiments, we fine-tuned the models on the source dataset with a fixed number of epochs and directly evaluated on the target dataset. We used the AdamW optimizer, with $\beta_{1}=0.9$, $\beta_{2}=0.95$, $eps=10^{-5}$. We fine-tuned the pretrained encoder and the classification head for 5 epochs with a learning rate of $1\times 10^{-5}$ and $1\times 10^{-4}$.
@@
SOZ localization. The SOZ localization was conducted on 4 patients with DRE implanted with 4 to 10 electrodes (47 to 120 channels). The data contain channel-level annotations. The sampling rate is 500 Hz and we split the data into 3-second segments. For a patient experienced $N$ seizures, BrainWave provided channel-level predicted probabilities $\mathbf{p}=\{\mathbf{p}_{i}^{c,t}\in[0,1]|i=1,2...,N;c=1,2,...,C;t=1,2,...,T_{i}\}$, where $C$ is the number of channels and $T_{i}$ is the number of segments during $i$’th seizure. For each channel, we calculated two metrics based on the predictions: the probability of epileptic discharge occurrence, and the number of times it serves as the seizure onset site. The first metric was obtained by a mean pooling of $\mathbf{p}$ along the channel axis: $\frac{1}{\sum_{i=1}^{N}T_{i}}\sum_{i=1}^{N}\sum_{t=1}^{T_{i}}\mathbf{p}_{i}^{c,t}\in\mathbb{R}^{C}$. The second metric was counted across $N$ seizures. For each seizure, we smoothed the probabilities with median filtering along the time axis. Then we identified the earliest channel(s) that generated epileptic waveforms, defined as: $c^{*}=\arg\min_{(c,t)\in\mathcal{S}_{i}}t$, where $\mathcal{S}_{i}$ is the set of $(c,t)$ pairs that satisfy $\text{MedianFilter}(\mathbf{p}_{i}^{c,t})\geq 0.5$. We counted the occurrences of each identified channel across $N$ seizures.
Prediction of clinical assessment scores in AD patients. The clinical scale scores we predicted in Sec. 2.3 included MMSE, MoCA-B, ROCF, and PSQI. We divided these scores into several discrete ranges. MMSE: 24-30, 21-23, 10-20 and 0-9. MoCA-B: 26-30, 18-25, 10-17 and 0-9. ROCF: 33-36, 24-32, 18-23, 12-17 and 0-11. PSQI: 0-5, 6-10, 11-15 and 16-21.
4.4 Data Availability
4.5 Code Availability
We will release the model weights, pretraining code, and usage code upon publication.